### v3

# FACT-GS: Frequency-Aligned Complexity-Aware Texture Reparameterization for 2D Gaussian Splatting

Abstract
Realistic scene appearance modeling has advanced rapidly with Gaussian Splatting, which enables real-time, high-quality rendering. Recent advances introduced per-primitive textures that incorporate spatial color variations within each Gaussian, improving their expressiveness. However, texture-based Gaussians parameterize appearance with a uniform per-Gaussian sampling grid, allocating equal sampling density regardless of local visual complexity, which leads to inefficient texture space utilization. We introduce FACT-GS, a Frequency-Aligned Complexity-Aware Texture Gaussian Splatting framework that allocates texture sampling density according to local visual frequency. Grounded in adaptive sampling theory, FACT-GS reformulates texture parameterization as a differentiable sampling-density allocation problem, replacing the uniform textures with a learnable frequency-aware allocation strategy implemented via a deformation field whose Jacobian modulates local sampling density. Built on 2D Gaussian Splatting, FACT-GS performs non-uniform sampling on fixed-resolution texture grids, preserving real-time performance while recovering sharper high-frequency details under the same parameter budget.
1 Introduction
Realistic 3D appearance modeling is essential for large-scale reconstruction, dynamic scene understanding, and photorealistic rendering [19, 23, 7], but achieving high fidelity under real-time constraints remains challenging. Implicit volumetric fields such as NeRF [25, 2, 42, 49] provide high visual quality but require dense sampling and costly MLP inference, limiting scalability. In contrast, 3D Gaussian Splatting (3DGS) [17, 46, 6] represents scenes using explicit differentiable Gaussian primitives, achieving real-time high-quality rendering. Building on this, 2D Gaussian Splatting [14] uses planar primitives that simplify ray–primitive interaction and improve geometric stability. However, 2D and 3D Gaussian splatting represent appearance using view-dependent spherical harmonics [17, 14], without spatial color variation within each primitive, constraining its expressiveness.
To address this limitation, texture-based Gaussian extensions [4, 28, 15] attach per-Gaussian spatially varying textures. However, the appearance parameterization of these methods is uniformly distributed across the texture space of each Gaussian, without adapting to local visual complexity [11, 37, 40]. This uniform allocation leads to inefficient utilization of the texture space: high-frequency regions (e.g., sharp edges, fine patterns, numerical markings) are allocated insufficient texture space, resulting in detail loss, while large smooth regions inefficiently consume capacity to represent nearly uniform color. In practice, each Gaussian can only store a low-resolution texture patch (e.g., $4\times 4\times 4$), exacerbating the expressiveness constraints imposed by uniform sampling. As shown in Fig. LABEL:fig:teaser, this results in blurred appearances in visually complex regions. Simply increasing texture resolution yields only marginal improvements at quadratically increased memory and bandwidth cost, while neural texture fields [44, 48] break the real-time rendering performance of the Gaussian-based representations.
Motivated by this structural inefficiency, we propose an adaptive sampling scheme of the per-Gaussian texture where sampling should follow the local signal frequency rather than remain spatially uniform. To achieve this, we introduce a frequency-aligned differentiable texture reparameterization that continuously modulates sampling density across the texture domain. Grounded in adaptive sampling theory [12, 29, 51], this allocates more sampling capacity to high-frequency regions and less to smooth areas. Building on this formulation, we present FACT-GS, a Frequency-Aligned Complexity-Aware Texture Gaussian Splatting framework that integrates this frequency-adaptive reparameterization.
Concretely, FACT-GS is built upon 2D Gaussian Splatting [14] and introduces a learnable deformation field defined in the texture domain. This field predicts continuous per-texel displacements, inducing a smooth warping whose Jacobian determinant [30] directly modulates local sampling density. In contrast to previous uniform texture parameterizations [4, 28, 15], which allocate equal resolution regardless of signal complexity, our warping adaptively concentrates samples in high-frequency regions while compressing low-frequency areas. This realizes adaptive sampling as a differentiable, per-primitive reparameterization, allowing texture capacity to naturally focus where visual detail is needed. The entire framework is trained end-to-end under standard photometric supervision, as the deformation field is designed to integrate seamlessly into the differentiable splatting pipeline. As a result, FACT-GS recovers sharper high-frequency appearance and uses parameters more efficiently, all while maintaining real-time rendering performance under the same primitive budget.
Our contributions are summarized as follows:
  • We propose FACT-GS, a frequency-aligned texture reparameterization that reallocates sampling density based on local signal frequency under a fixed texture budget.
  • We reformulate the texturing problem through the lens of adaptive sampling theory as a sampling-density allocation problem, achieved via a learnable deformation field whose Jacobian controls local texture space distortion.
  • We demonstrate through extensive qualitative, quantitative experiments, and ablation studies that our method provides consistent, detail-preserving improvements over state-of-the-art textured Gaussian methods across benchmarks without sacrificing real-time performance.
2 Related Work
Novel View Synthesis. Novel view synthesis (NVS) aims to generate photorealistic renderings of 3D scenes from unseen viewpoints, typically given multi-view images with known camera poses. Its development originates from classical SfM and MVS pipelines [33, 34, 1, 31, 32, 10, 9], which reconstruct geometry through multi-view correspondences but suffer from occlusion sensitivity and limited viewpoint extrapolation. Neural radiance fields (NeRF) [25] later revolutionized NVS by representing scenes as implicit neural functions optimized via photometric supervision. Subsequent extensions improve anti-aliasing [2], large-scale reconstruction [38, 3], reflection modeling [39], and acceleration via grid-based encodings [49], yet they remain constrained by dense ray sampling and slow MLP inference.
This motivates explicit representations such as 3D Gaussian Splatting [17], which replace neural volumes with differentiable Gaussian primitives for real-time rendering. Building upon this foundation, follow-up works improve geometric accuracy and rendering stability through enhanced kernel formulations and densification strategies [8, 20, 21]. To further refine geometry, 2D Gaussian Splatting [14] introduces planar Gaussian primitives for more accurate and efficient radiance field representation, enabling fine-grained geometric reconstruction while maintaining high appearance quality. However, both 3D and 2D Gaussian Splatting modeled appearance via view-dependent spherical harmonic coefficients, but lack spatial color variation within each primitive, limiting their expressiveness.
Spatially Varying Gaussians. Recent works introduce spatially varying Gaussians to overcome this constraint. One line of research augments each primitive with RGB or RGBA textures to model spatially varying colors [4, 43, 28, 15, 35] or surface normals for relighting [36]. Another employs neural texture fields to decouple geometry and appearance [44, 48], achieving high visual fidelity at the cost of slower training and rendering due to MLP inference.
Texture-based Gaussians strike a balance between quality and efficiency, providing stronger appearance modeling without sacrificing real-time performance. However, their uniformly sampled texture parameterization remains inefficient, failing to allocate texture capacity according to signal complexity. This uniform sampling strategy leads to a structural sampling–complexity mismatch, where high-frequency regions receive insufficient capacity while smooth regions waste parameters.
Adaptive Sampling and Frequency-Aware Parameterization. Classical rendering and Monte Carlo integration extensively explored adaptive sampling [11, 51, 29, 37, 40], where sample density is modulated by local signal frequency or reconstruction error to achieve efficient high-frequency reconstruction. At the representation level, multi-resolution texture filtering such as mipmapping and anisotropic filtering [13, 41, 27] pre-filters textures across scales to suppress aliasing and maintain visual consistency under varying viewing distances. Recent neural and Gaussian-based formulations [22, 50, 24] extend this principle to learned 3D representations, enforcing frequency- or scale-consistent parameterization in both geometry and appearance reconstruction. However, these methods operate at the global or feature-map level and still assume uniform sampling within each primitive, limiting their ability to adapt texture-space resolution to local signal variation.
To address this, we introduce a differentiable frequency-aligned reparameterization that operates at the per-primitive coordinate level, bridging adaptive sampling theory with texture parameterization in Gaussian representations.
3 Preliminaries
Figure 2: Method comparison of Textured GS and FACT-GS. (A) For the $i$-th Gaussian, the ray–Gaussian intersection was firstly computed to obtain texture coordinates $(u,v)$. Textured GS bilinearly sampled from RGB and opacity textures $\mathbf{T}_{i}^{\text{RGB}}$ and $\mathbf{T}_{i}^{\alpha}$ (B) using $(u,v)$ to get the texture color $c_{i}^{\text{tex}}$ and opacity $\alpha_{i}^{\text{tex}}$ (C), which lead to blurred high-frequency details (edges). Instead, FACT-GS uses a learnable deformation field $\mathbf{D}_{i}$ to predict a continuous warp $\Phi_{i}(u,v)$ that allocate sampling density guided by local frequency (D), after which the warped coordinates $\Phi_{i}(u,v)$ are used to sample from RGB and opacity textures (E), producing the FACT texture color and opacity (F), which has more clear high-frequency details.
3.1 2D Gaussian Splatting Representation
We adopt the 2D Gaussian Splatting representation [14]. Each primitive lies on a local tangent plane with center $p_{k}$, orthonormal basis $(t_{\beta},t_{\gamma},n_{k})$, and scales $(s_{\beta},s_{\gamma})$. Local coordinates on the plane are $(\beta,\gamma)$:
$P(\beta,\gamma)=p_{k}+s_{\beta}t_{\beta}\beta+s_{\gamma}t_{\gamma}\gamma.$ (1)
The spatial weight is $\mathcal{G}(\beta,\gamma)=\exp[-(\beta^{2}+\gamma^{2})/2]$, and splats are projected and composited in image space. Appearance is modeled using spherical harmonics, producing one view-dependent color per primitive, without spatial variation across the tangent plane.
3.2 Textured Gaussians
Standard 2D Gaussians provide only a single view-dependent color for each primitive. Textured Gaussians [4] instead attach a learnable RGBA texture map $\mathbf{T}_{i}\in\mathbb{R}^{\tau\times\tau\times 4}$ to each primitive, enabling spatially varying appearance within tangent plane. Although originally formulated in 3D, the official implementation adopts a 2D variant for more accurate and efficient ray–Gaussian intersection, which we also follow.
Given the local intersection coordinates $\mathbf{s}(\mathbf{x})=(\beta,\gamma)$ on the tangent plane, the corresponding texture coordinates $(u,v)$ are computed as
$u=\frac{\beta+\xi}{2\xi}\tau,\quad v=\frac{\gamma+\xi}{2\xi}\tau,$ (2)
where $\xi$ controls the effective texture support. The texture color $c_{i}^{\mathrm{tex}}(\mathbf{p})$ and opacity $\alpha_{i}^{\mathrm{tex}}(\mathbf{p})$ are bilinearly sampled from the texture map $\mathbf{T}_{i}(u,v)$. We combine these with the SH-encoded base color $c_{i}^{\mathrm{base}}(\mathbf{p})$ and base opacity $o_{i}$, and composite the final color using volumetric alpha blending weighted by the Gaussian footprint $\mathcal{G}_{i}(\mathbf{s}(\mathbf{x}))$:
$c(\mathbf{p})$ $=\sum_{i=1}^{N}\bigl(c_{i}^{\mathrm{tex}}(\mathbf{p})+c_{i}^{\mathrm{base}}(\mathbf{p})\bigr)\,\bigl(\alpha_{i}^{\mathrm{tex}}(\mathbf{p})\,\mathcal{G}_{i}(\mathbf{s}(\mathbf{x}))\,o_{i}\bigr)$ (3)
$\quad\times\prod_{j<i}\Bigl[1-\alpha_{j}^{\mathrm{tex}}(\mathbf{p})\,\mathcal{G}_{j}(\mathbf{s}(\mathbf{x}))\,o_{j}\Bigr].$
4 Method
4.1 Motivation
While Textured Gaussians introduce spatial appearance variation through learnable texture maps $\mathbf{T}_{i}$, these textures are parameterized on a uniform grid that allocates equal representational capacity within each splat regardless of local signal frequency. As a result, high-frequency regions that occupy small spatial areas are under-represented, whereas smooth regions that span larger areas overconsume parameters to represent redundant information, leading to a persistent sampling–complexity mismatch in the texture domain and imbalanced texture capacity allocation.
This mismatch manifests in two characteristic failure modes: (1) Detail loss. High-frequency regions are assigned too few texels, resulting in insufficient representational capacity to capture fine details (see Fig. LABEL:fig:teaser; quantified in Fig. 5). (2) Inefficient resolution scaling. Increasing texture resolution raises memory and bandwidth quadratically, yet provides only marginal visual improvement since the sampling pattern remains uniform.
Fundamentally, this imbalance reflects a disconnect between texture capacity and the spatial frequency of the underlying appearance, motivating the need for frequency-aware allocation. To address the sampling–complexity mismatch in texture space, we reinterpret texture parameterization as a sampling-density allocation problem through the lens of adaptive sampling theory.
Let $C(u,v)$ denote the ground-truth color function, and let the texture $\mathbf{T}$ be its discrete representation, where texels sample $C(u,v)$ in texture space. The sampling density $\rho(u,v)$ represents the number of texels per unit area. Uniform parameterization is structurally inefficient because it assigns the same density everywhere despite large spatial variation in the underlying appearance signal [30, 51].
Classical reconstruction theory suggests that sampling density should increase in regions where the signal varies more rapidly. Since a texture map is a discrete sampling of the continuous appearance function $C(u,v)$, its local variation behaves analogously to 2D image signals in reconstruction theory: higher spatial variation corresponds to higher local frequency and therefore requires denser sampling.
Following prior work that uses local variation as a proxy for frequency content [5, 29], we adopt the gradient-based target density
$\rho^{\star}(u,v)\propto(\|\nabla C(u,v)\|+\epsilon)^{\alpha},$ (4)
where $\|\nabla C\|$ estimates local spatial frequency, $\alpha$ controls adaptation strength, and $\epsilon$ ensures stability.
Under a continuous reparameterization $\Phi$, a region around $(u,v)$ in the parameter space with area $dA$ is mapped to a region with area $|\det J_{\Phi}(u,v)|\,dA$ in warped texture space, therefore the induced sampling density is modulated by $|\det J_{\Phi}(u,v)|$. Thus, a larger $|\det J_{\Phi}|$ indicates higher local texture capacity. Matching $\rho^{\star}(u,v)$ therefore amounts to shaping the Jacobian determinant of $\Phi$. Because photometric loss concentrates gradients in high-frequency regions, optimization naturally expands $\Phi$ in these areas, increasing $|\det J_{\Phi}|$ until the induced density aligns with $\rho^{\star}$.
These observations motivate treating texture space as a deformable sampling domain whose density adapts to local visual complexity. We next introduce a frequency-aligned reparameterization that realizes this idea.
4.2 Frequency-Aligned Complexity-Aware Texture Reparameterization
We realize the target sampling density $\rho^{\star}(u,v)$ by introducing a learnable warp $\Phi:(u,v)\mapsto(u^{\prime},v^{\prime})$ over texture space. By reshaping the domain, the warp redistributes texels according to signal complexity: high-frequency regions are expanded to receive more texels, while smooth regions are compressed to avoid redundant sampling. This redistribution is realized through the local area change of the warp, captured by the Jacobian determinant $|\det J_{\Phi}(u,v)|$.
In contrast to texture-based Gaussians [4, 28], which operate on fixed uniform grids and primarily improve detail by increasing resolution, our formulation adjusts sampling density geometrically through $\Phi$ while keeping the grid resolution fixed. This enables non-uniform, complexity-aware texture parameterization that remains fully differentiable and end-to-end trainable.
(i) Differentiable Warping Field. For each Gaussian, we instantiate $\Phi$ using a deformation field $\mathbf{D}_{i}\in\mathbb{R}^{\tau\times\tau\times 2}$ that predicts per-texel displacements, $\mathbf{D}_{i}(u,v)=(D_{i}^{u}(u,v),D_{i}^{v}(u,v))$. Warped coordinates are obtained as
$(u^{\prime},v^{\prime})=(u,v)+\lambda\,\mathbf{D}_{i}(u,v),$ (5)
defining the continuous mapping $\Phi_{i}$. The Jacobian $J_{\Phi_{i}}$ captures the resulting local area distortion and therefore the effective sampling density at each texel.
(ii) Texture Sampling and Gradient Modulation. The warped coordinates $\Phi_{i}(u,v)$ are used for differentiable texture lookup:
$c_{i}^{\mathrm{tex}}(\mathbf{p})=\mathbf{T}_{i}^{\mathrm{RGB}}\!\big(\Phi_{i}(u,v)\big),\quad\alpha_{i}^{\mathrm{tex}}(\mathbf{p})=\mathbf{T}_{i}^{\alpha}\!\big(\Phi_{i}(u,v)\big).$ (6)
Differentiating Eq. (6) with respect to $(u,v)$ and applying the chain rule yields:
$\nabla c_{i}^{\mathrm{tex}}(u,v)=J_{\Phi_{i}}(u,v)^{\!\top}\,\nabla\mathbf{T}_{i}^{\mathrm{RGB}}\!\big(\Phi_{i}(u,v)\big).$ (7)
Here, $\nabla\mathbf{T}_{i}^{\mathrm{RGB}}$ denotes the spatial gradient of the uniform-grid texture. This expression shows that the warp reshapes the texture’s local gradient field through $J_{\Phi_{i}}^{\top}$, effectively modulating local frequency content.
The local signal frequency can be approximated by the gradient magnitude $\|\nabla c_{i}^{\mathrm{tex}}(u,v)\|$ [5]. Under the warp, the uniform-grid texture gradient $\nabla\mathbf{T}_{i}^{\mathrm{RGB}}$ is transformed by the singular values of $J_{\Phi_{i}}$, which respectively amplify or attenuate its magnitude. These effects manifest as regions receiving more texels allocated or fewer texels received, enabling the FACT texture $c_{i}^{\mathrm{tex}}(u,v)$ to respond faithfully to higher-frequency variation in the ground-truth color function $C(u,v)$.
This establishes the connection between geometric reparameterization and frequency-aligned density allocation. The deformation field $\mathbf{D}_{i}$ is trained end-to-end via the photometric loss, which implicitly drives $|\det J_{\Phi_{i}}(u,v)|$ toward the target density $\rho^{\star}(u,v)$.
(iii) Final Rendering. After warping, the per-Gaussian colors and opacities are evaluated at $\Phi_{i}(u,v)$ and composited using the standard differentiable splatting procedure in Eq. (3). The rendering pipeline itself remains unchanged; only the texture parameterization is modified, ensuring that real-time Gaussian splatting is fully preserved.
The complete FACT-GS pipeline is illustrated in Fig. 2, and Alg. 1 summarizes the training process.
Input: Pretrained Gaussian parameters $\{p_{i},s_{\beta},s_{\gamma},t_{\beta},t_{\gamma},o_{i},\mathrm{SH}_{i}\}$, textures $\{\mathbf{T}_{i}\}$ and deformation field $\{\mathbf{D}_{i}\}$
Output: Rendered image $\hat{I}$ with aligned details
1
2for each Gaussian $i=1,\dots,N$ do
3   Compute $(u,v)$ via tangent-plane mapping;
4   Predict displacement $\mathbf{D}_{i}(u,v)$;
5   Warp coordinates $(u^{\prime},v^{\prime})=(u,v)+\mathbf{D}_{i}(u,v)$;
6   Sample $c_{i}^{\mathrm{tex}},\alpha_{i}^{\mathrm{tex}}$ from $\mathbf{T}_{i}(u^{\prime},v^{\prime})$;
7
8
9Blend all Gaussians via Eq. (3) to get $\hat{I}$;
10 Backpropagate loss to update all parameters.
Algorithm 1 Training of FACT-GS
Figure 3: Novel view synthesis comparison under reduced primitive budgets (10% / 1% of default 2DGS, exact budgets are shown in Supp. 7.3). 2DGS [14] loses high-frequency appearance, and Textured GS [4] blurs fine detail due to uniform texture sampling. Our method preserves sharp edges, textures, and material details (zoom-ins), achieving significantly higher perceptual fidelity without increasing parameters.
Discussion. Our formulation reframes texture parameterization as an adaptive sampling problem: the objective is no longer to uniformly cover the texture domain, but to allocate representational capacity in proportion to local signal frequency. The continuous mapping $\Phi$ provides a learnable mechanism that moves the effective sampling density toward the frequency-aligned target, shifting texture layouts from geometry-driven to information-driven organization.
Although instantiated on Gaussian textures, the principle is general: any spatially parameterized appearance representation (e.g., feature planes, neural textures, volumetric grids) can adopt frequency-aligned sampling to improve the distribution of representational capacity.
In essence, FACT-GS replaces uniform parameterization with a dynamic, frequency-aware allocation rule, enabling higher visual fidelity under the same parameter budget.
4.3 Optimization
Following Textured Gaussians [4], we adopt a two-stage optimization for stability. In the first stage, the 2D Gaussian primitives are trained using the standard losses from 3D Gaussian Splatting [17], combining $\mathcal{L}_{1}$ photometric and structural similarity ($\mathcal{L}_{\text{SSIM}}$) terms, augmented with a mask loss $\mathcal{L}_{\alpha}$ when ground-truth masks are available:
$\mathcal{L}=\eta\mathcal{L}_{1}+(1-\eta)\mathcal{L}_{\text{SSIM}}+\mathcal{L}_{\alpha},\quad\eta=0.2.$ (8)
This stage runs for 30,00030{,}000 iterations. In the second stage, we optimize the per-Gaussian texture $\mathbf{T}$ and deformation field $\mathbf{D}$ while fine-tuning the Gaussian parameters $\{p_{k},s_{\beta},s_{\gamma},t_{\beta},t_{\gamma},o,\text{SH}\}$ without pruning or densification, for another 30,00030{,}000 iterations. Learning rates for $\mathbf{T}$ and $\mathbf{D}$ are set to $2.5\times 10^{-3}$ and $1\times 10^{-3}$, respectively.
5 Experiments
5.1 Implementation Details
We build on the gsplat [45] framework and extend it with a lightweight CUDA kernel for frequency-aligned texture warping. The per-Gaussian texture space sampling-density modulation described in section 4.2 introduces negligible overhead during inference. All experiments are run on a single NVIDIA RTX A6000 GPU.
5.2 Datasets and Evaluation Metrics
Dataset. We evaluate our method on five standard benchmarks for novel view synthesis: NeRF Synthetic [25] (8 scenes), MipNeRF 360 v2 [3] (7 scenes), DTU [16] (4 scenes), Tanks & Temples [18] (3 scenes), and LLFF [26] (2 scenes).
Metrics. Following Textured GS [4], we report Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [47] for quantitative evaluation.
5.3 Experimental Results
Methods NeRF Synthetic [25] MipNeRF 360v2 [3] DTU [16] Tanks & Temples [18] LLFF [26]
[…]
2D GS (100%) 33.38 0.966 0.0275 28.96 0.870 0.0997 27.85 0.905 0.1328 22.79 0.823 0.1572 27.32 0.883 0.1092
Textured GS (100%) 33.91 0.968 0.0235 29.30 0.873 0.0915 28.76 0.905 0.1241 23.56 0.831 0.1428 29.04 0.899 0.0852
Ours (100%) 34.02 0.969 0.0220 29.34 0.874 0.0889 28.76 0.906 0.1204 23.53 0.834 0.1377 28.99 0.900 0.0827
2D GS (10%) 30.35 0.945 0.0594 27.22 0.805 0.1937 29.04 0.899 0.1791 22.18 0.767 0.2577 27.37 0.863 0.1491
Textured GS (10%) 30.88 0.949 0.0542 27.80 0.825 0.1646 28.72 0.894 0.1640 22.65 0.792 0.2206 28.29 0.884 0.1206
Ours (10%) 31.51 0.954 0.0439 28.06 0.837 0.1447 28.79 0.904 0.1504 22.83 0.802 0.2036 28.46 0.891 0.1058
2D GS (1%) 25.01 0.885 0.1621 23.76 0.648 0.4172 26.99 0.847 0.2963 19.68 0.644 0.4416 24.83 0.766 0.3041
Textured GS (1%) 25.54 0.892 0.1523 24.38 0.675 0.3710 27.02 0.851 0.2724 20.25 0.677 0.3954 25.23 0.788 0.2763
Ours (1%) 26.44 0.904 0.1191 24.84 0.703 0.3247 27.46 0.865 0.2417 20.60 0.698 0.3579 26.00 0.818 0.2291
Table 1: Quantitative comparison with 2DGS [14] and Textured GS [4] under varying numbers of Gaussian primitives. Our method consistently achieves higher fidelity across different primitive budgets. (Best scores are shown in bold, second-best are underlined.)
Quantitative Results. We evaluate our method against Textured GS [4] and 2D GS [14] under varying primitive budgets ($0\%, 1010\%$, $1\%$). All methods follow the same 2D GS training protocol [14]. Table 1 summarizes the results; corresponding average primitive counts are provided in the appendix. To ensure comparable parameter budgets, we match per-Gaussian texture capacity by setting $\tau_{\text{tex}}=5$ for Textured GS and $\tau_{\text{FACT}}=4$ for our method.
As shown in Table 1, our method achieves consistently higher PSNR and SSIM, and lower LPIPS across all datasets. The advantage becomes most pronounced under low primitive budgets ($10\%$ and $1\%$), where FACT-GS improves PSNR by up to +1.0 dB and reduces LPIPS by 22%. This behavior is expected: when representational capacity is limited, uniform texture sampling wastes parameters in smooth regions, whereas our frequency-aligned parameterization concentrates capacity on high-frequency structures.
We further evaluate our method ($100\%$ primitive budget) on the NeRF Synthetic [25] and MipNeRF 360-Indoor [3] datasets. We compare against Textured GS [4], SuperGS [43], GsTex [28], and Neural Shell [48].
As shown in Table 2, our method achieves the highest overall fidelity while maintaining real-time rendering performance (in Table 3). In contrast, methods such as Neural Shell [48] improve accuracy only by increasing computation. This indicates that our method does not rely on capacity scaling, but instead reallocates texture resolution to high-frequency regions—yielding structural efficiency gains directly predicted by adaptive sampling theory.
Detail Preservation Under Limited Primitive Budgets. Fig. 3 compares novel view synthesis when using only $10\%$ and $1\%$ of the default 2DGS primitive count. 2D GS [14] loses fine textures entirely due to its single-color appearance model, and Textured GS [4] produces blurred patterns because its uniform texture grid allocates capacity evenly across the surface. In contrast, our method preserves sharp edges, printed characters, and high-frequency materials by reallocating sampling density to regions that require detail. This confirms that frequency-aligned parameterization maintains perceptual fidelity even when representational resources are severely constrained.
Detail Preservation in High-Capacity Regime. Fig. 4 compares reconstruction under high primitive budgets ($100\%$ for Lego, $50\%$ for Church). Even with many primitives, 2D GS loses fine texture entirely due to its single-color primitive model, while Textured GS still blurs high-frequency patterns because its uniform texture grid allocates texels evenly, independent of local detail.
In contrast, FACT-GS preserves thin structures, decorative patterns, and high-frequency materials. The improvement is structural: frequency-aligned parameterization concentrates sampling density in visually complex regions instead of wasting capacity in smooth areas.
Thus, our method improves detail fidelity even when the primitive budget is not constrained, demonstrating that the gain arises from better sampling allocation, not from increased capacity.
Methods NeRF Synthetic [25] MipNeRF 360-indoor [3]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS [17] 33.34 0.969 0.030 31.03 0.921 0.188
SuperGS [43] 33.71 0.970 0.031 30.23 0.917 0.188
GsTex [28] 33.37 0.965 0.041 30.46 0.915 0.204
Neural Shell [48] 33.50 0.967 0.032 30.59 0.911 0.174
Textured GS [4] (100%) 33.91 0.968 0.024 31.71 0.938 0.061
Ours(100%) 34.02 0.969 0.022 31.80 0.940 0.059
Table 2: Quantitative comparison with prior Gaussian-based methods, including 3DGS [17], Neural Shell [48], SuperGS [43], GsTex [28], and Textured GS [4]. All methods are evaluated using the full (100%) 2DGS primitive budget [14] on NeRF Synthetic [25] and MipNeRF 360-Indoor [3]. (Best results are shown in bold, and second-best are underlined.)
Figure 4: Qualitative comparison under high primitive budgets ($100\%$ for Lego, $50\%$ for Church). 2D GS [14] lacks spatial texture variation, and Textured GS [4] shows blurred high-frequency patterns due to uniform texture sampling. Our method preserves thin structures and fine texture details.
Inference Speed (FPS) Textured GS [15] Neural Shell [48] Ours
NeRF Synthetic (100%) 121 71 119
NeRF Synthetic (10%) 216 - 215
[…]
5.4 Runtime and Efficiency
As shown in Table 3, FACT-GS maintains real-time rendering and remains within $1\text{–}2$ FPS of Textured GS [4] on NeRF Synthetic [25] using a single RTX A6000 GPU. The sampling-density alignment is implemented as a per-primitive CUDA warp computed in parallel, keeping the rasterization pipeline unchanged and thus introducing negligible overhead.
5.5 Ablation and Analysis
Figure 5: Per-Gaussian average RGB gradient magnitude. Textured GS [15] concentrates most textures in the low-frequency range (<10<10), indicating insufficient capacity for high-frequency regions. Our method reallocates capacity according to local signal complexity, yielding more high-frequency textures under the same parameter budget.
Texture Frequency Analysis. To quantify the effect of our reparameterization, we measure the spatial frequency of each learned per-Gaussian RGB texture $\mathbf{T}\in\mathbb{R}^{\tau\times\tau\times 3}$ using the mean gradient magnitude [5]:
$\text{Freq}=\frac{1}{3\tau^{2}}\sum_{c\in\{R,G,B\}}\sum_{\mathbf{p}\in\mathbf{T}}\|g(\mathbf{p}_{c})\|,$ (9)
where $g(\cdot)$ is the Sobel gradient.
Fig. 5 shows the resulting distributions. Visually, Textured GS [15] forms a narrow peak in the low-frequency range (<10<10), indicating that its uniform grid under-samples visually complex regions. In contrast, FACT-GS produces a broader and right-shifted distribution, reflecting substantially more high-frequency textures under the same parameter budget.
This shows that FACT-GS does not merely sharpen appearance; it reallocates sampling capacity toward high-frequency regions, aligning texture resolution with signal complexity. In other words, the observed gains arise not from additional parameters, but from frequency-aligned sampling, as predicted by our theoretical formulation.
Ablation on Texture Parameters. Fig. 6 evaluates reconstruction quality as the number of per-Gaussian texture parameters increases from $100$ to $3600$, using $10\%$ of the default 2DGS primitives. Across all parameter budgets, FACT-GS achieves higher PSNR than Textured GS [15]. Notably, FACT-GS matches the reconstruction quality of Textured GS while using only $\tfrac{1}{16}$ of its texture parameters. These results confirm that frequency-aligned sampling enhances the efficiency of texture space utilization compared to uniform-grid textures, offering substantially greater improvements than simply increasing the texture resolution.
Figure 6: Ablation on the number of per-Gaussian texture parameters (PSNR, $10\%$ primitive budget). As texture capacity increases, FACT-GS maintains the lead across all parameter settings. Our method achieves the same performance using only $\tfrac{1}{16}$ of the texture parameters.
6 Conclusion
We present FACT-GS, a frequency-aligned texture Gaussian framework that reformulates texture parameterization as an adaptive sampling problem. By coupling sampling density with local signal frequency through a learnable deformation field, FACT-GS enables adaptive, non-uniform sampling, yielding sharper details and significantly higher texture efficiency while preserving real-time rendering performance under the same parameter budget.
More broadly, this work shifts appearance modeling from uniform parameterization to signal-driven adaptive parameterization: assigning capacity according to information density rather than uniformly, which provides a foundation for future neural rendering methods guided by signal complexity.
Limitations and Future Work. FACT-GS currently only looks at the gradient of the texture and focuses on spatial alignment at the per-texture level. In the future, we are looking at incorporating explicit frequency predictors and exploring the extension of this technique to dynamic scenes.
Acknowledgment
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), under funding reference numbers RGPIN-2021-03477.
Supplementary Material
7 Appendix
7.1 Author Contributions
Tianhao Xie and Linlian Jiang contributed equally to this work. Tianhao Xie designed the algorithm, implemented it, and conducted the experiments. Linlian Jiang led the method formulation and was primarily responsible for manuscript writing.
7.2 Datasets
We evaluate our method on five standard benchmarks for novel view synthesis, covering both synthetic and real-world scenes. The experimental scenes are listed below:
  • NeRF Synthetic [25]: all scenes.
  • MipNeRF 360 v2 [3]: publicly available scenes including Bicycle, Counter, Garden, Room, Bonsai, Kitchen, and Stump.
  • DTU [16]: Scan105, Scan110, Scan37, and Scan63.
  • Tanks & Temples [18]: Train, Truck, and Church.
  • LLFF [26]: Horns and Fortress.
7.3 Budgets for Fig. 3 and Fig. LABEL:fig:teaser
  • NeRF Ship: $10\%$.
  • NeRF Chair: $10\%$
  • NeRF Hotdog: $1\%$
  • Tanks & Temples Truck: $1\%$.
  • Tanks & Temples Train: $1\%$
  • Tanks & Temples Church: $10\%$
  • MipNeRF 360 Kitchen: $10\%$.
7.4 Ablation on Texture Parameters
[…]
Across all parameter budgets and datasets, FACT-GS achieves higher PSNR and SSIM, and lower LPIPS than Textured GS [15], which demonstrates that our method improves the efficiency of texture space utilization compared with uniform-grid textures, yielding substantially greater performance gains than simply increasing the texture resolution across datasets.
Parameters Count. To ensure a fair comparison with Textured GS, we matched the parameter number of both methods as closely as possible; nevertheless, our method consistently uses fewer parameters. The texture parameters of Textured GS are computed as $\tau_{\text{tex}}\times\tau_{\text{tex}}\times 4$, while those of our method are $\tau_{\text{FACT}}\times\tau_{\text{FACT}}\times 6$. We report the exact number of texture parameters for both methods in Table 4.
7.5 Model Size
Model size is matched for TexturedGS and our method in the main paper with identical Gaussian counts and total texture parameter budgets; an explicit size-matched comparison with 2DGS is further reported in Tab. 5 and Fig. 9.
7.6 Interpreting RGB Gradient.
We provide additional qualitative evidence for texture capacity utilization in Fig. 7. RGB gradient magnitude provides a practical proxy for local spatial frequency and effective texel allocation under a fixed texel budget, as higher gradients correspond to more rapidly varying color signals. Fig. 7(C–D) visualizes per-Gaussian local texture footprints, where our method concentrates color variation into a more compact and structured support, producing sharper transitions and clearer edges than TexturedGS. Accordingly, Fig.5 shows a right-shift in the RGB gradient distribution under parameter-matched settings, indicating increased effective utilization of texture capacity for high-frequency detail reconstruction.
Figure 7: RGB Gradient and per-Gaussian Texture Visualization.
[…]
Textured GS Ours
$5\times 5\times 4=100$ $4\times 4\times 6=96$
$10\times 10\times 4=400$ $8\times 8\times 6=384$
$15\times 15\times 4=900$ $12\times 12\times 6=864$
$20\times 20\times 4=1600$ $16\times 16\times 6=1536$
$30\times 30\times 4=3600$ $24\times 24\times 6=3456$
Table 4: Exact texture parameters count of Textured GS and our method for the ablation results in Fig. 8.
Number of Primitives. We report the average number of primitives in the default 100% 2DGS setting in Table 6.
Method Average Size Tanks &Temples LLFF DTU
(MB) PSNR ($\uparrow$)  /  SSIM ($\uparrow$)  /  LPIPS ($\downarrow$)
2DGS 100.94 22.79 / 0.805 / 0.194 27.68 / 0.881 / 0.113 27.85 / 0.905 / 0.133
Ours 99.23 23.11 / 0.816 / 0.177 28.46 / 0.891 / 0.106 28.90 / 0.910 / 0.128
Table 5: Size-matched comparison with 2DGS.
Figure 9: Quantitative size-matched comparison with 2DGS.
7.7 Backpropagation of FACT Texture
Forward. Let $\mathbf{T}_{i}\in\mathbb{R}^{\tau\times\tau\times 4}$ denote the texture map of the $i$-th Gaussian, and $\mathbf{D}_{i}\in\mathbb{R}^{\tau\times\tau\times 2}$ its learned deformation field. Given a camera ray from pixel $\mathbf{p}$ that intersects the $i$-th Gaussian at local coordinates $(u,v)$, the sampled displacement $(\Delta u,\Delta v)$ is computed by bilinear interpolation over $\mathbf{D}_{i}$:
$\Delta u=\sum_{j\in N_{(u,v)}}b_{j}(u,v)\,D_{j}^{u},\hskip 28.80008pt\Delta v=\sum_{j\in N_{(u,v)}}b_{j}(u,v)\,D_{j}^{v},$ (10)
where $N_{(u,v)}$ denotes the four neighboring texels surrounding $(u,v)$, and $b_{j}(u,v)$ are their corresponding bilinear interpolation weights. Let $\Phi(u,v)=(u,v)+(\Delta u,\Delta v)$ be the frequency-aligned sampling coordinate. The sampled color $c$ from texture $\mathbf{T}_{i}$ is then obtained as
$c=\sum_{k\in N_{\Phi(u,v)}}b_{k}(\Phi(u,v))\,T_{k},$ (11)
where $N_{\Phi(u,v)}$ are the four neighboring texels around $\Phi(u,v)$ and $b_{k}(\Phi(u,v))$ are the corresponding interpolation weights.
Backward. The gradient of the sampled color $c$ with respect to $D_{j}^{u}$ at texel $j$ is given by
$\frac{\partial c}{\partial D_{j}^{u}}=\Big(\sum_{k\in N_{\Phi(u,v)}}T_{k}\,\frac{\partial b_{k}(\Phi(u,v))}{\partial(u+\Delta u)}\Big)\frac{\partial\Delta u}{\partial D_{j}^{u}}.$ (12)
Substituting Eq. (10) into Eq. (12) yields
$\frac{\partial c}{\partial D_{j}^{u}}=\Big(\sum_{k\in N_{\Phi(u,v)}}T_{k}\,\frac{\partial b_{k}(\Phi(u,v))}{\partial(u+\Delta u)}\Big)b_{j}(u,v),$ (13)
which defines the differentiable gradient flow from the rendered color $c$ to the deformation field $\mathbf{D}_{i}$.
Number of primitives
NeRF Synthetics [25] 91,396
MipNeRF 360 v2 [3] 2,763,229
DTU [16] 316,989
Tanks & Temple [18] 1,584,660
LLFF [26] 490,511
Table 6: The average number of primitives under the default 100% optimization setting.
Methods NeRF Synthetic [25] MipNeRF 360v2 [3] DTU [16] Tanks & Temples [18] LLFF [26]
[…]
Textured GS (100) 30.88 0.949 0.0542 24.38 0.675 0.3710 28.72 0.899 0.1640 22.65 0.792 0.2206 28.29 0.884 0.1206
Ours (96) 31.51 0.954 0.0439 24.84 0.703 0.3247 28.80 0.904 0.1504 22.83 0.802 0.2036 28.46 0.891 0.1058
Textured GS (400) 31.33 0.953 0.0457 24.72 0.695 0.3242 28.46 0.898 0.1543 22.80 0.801 0.1974 28.36 0.885 0.1084
Ours (384) 31.89 0.957 0.0364 25.25 0.728 0.2710 28.51 0.901 0.1431 23.00 0.810 0.1791 28.68 0.896 0.0952
Textured GS (900) 31.49 0.954 0.0411 24.82 0.702 0.3000 28.31 0.898 0.1513 22.84 0.803 0.1847 28.17 0.886 0.1075
Ours (864) 31.90 0.957 0.0341 25.31 0.734 0.2506 28.61 0.901 0.1382 23.05 0.813 0.1684 28.60 0.894 0.0932
Textured GS (1600) 31.55 0.954 0.0383 24.85 0.707 0.2850 28.39 0.897 0.1475 22.86 0.805 0.1760 28.09 0.886 0.1037
Ours (1536) 31.87 0.957 0.0331 25.31 0.733 0.2420 28.48 0.900 0.1409 23.06 0.814 0.1619 28.49 0.894 0.0900
Textured GS (3600) 31.52 0.954 0.0359 24.82 0.707 0.2705 27.26 0.897 0.1441 22.84 0.804 0.1669 27.85 0.883 0.1013
Ours (3456) 31.73 0.956 0.0326 25.25 0.731 0.2374 28.44 0.898 0.1383 22.96 0.811 0.1567 28.32 0.889 0.0960
[…]